calculate the percentage of nitrogen content by mass in urea, CO(NH2)2.


1
Expert's answer
2021-05-24T07:19:02-0400

percentage of nitrogen content by mass in urea is given by ="\dfrac{\ mass\ of \ nitrogen}{molar\ mass \ of\ urea}\times 100"



Molecular formula of urea is given by = "NH_{2}-CO-NH_{2}"



Molar mass of urea = (14+1"\times 2)\times 2 +12 + 16=60 \ g/mol"



Mass of nitrogen in urea ="14\ g/mol \times 2 =28\ g/mol"


% of nitrogen content by mass in urea = "\dfrac{28\ g/mol}{60\ g/mol }\times 100" =46.6%